I have two questions:-
1>A bit is at a distance L from one end of a long rod.Coefficient of friction between the rod and the bit is “meu”. The rod is moving with uniform accelaration ”alfa” about the vertical axis.The rod is always horizontal.After how long the bit will start sliding?There is no gravity.
2>A body is moving with constant velocity in a straight line parallel axis.Can he has any variable or constant angular momentum?Explain please!!!!

Grade:11

2 Answers

Shobhit Varshney IIT Roorkee
9 years ago
dear student,

Question-1
At t=0 , angular velocity of rod (w) = 0
At t, bead starts to slide,then, wt = \alphat

Let coeff of friction = f, then f * (mg) = m wt2 L
Get the value of f.

Question-2
Yes he can constant and variable angular momentum depending upon the point or axix about which we are calculating the angular momentum.

Thanks.
Shobhit Varshney IIT Roorkee
9 years ago
dear student,

Question-1
At t=0 , angular velocity of rod (w) = 0
At t, bead starts to slide,then, wt = [\alpha] t

Let coeff of friction = f, then f * (mg) = m wt2 L
Get the value of f.

Question-2
Yes he can constant and variable angular momentum depending upon the point or axix about which we are calculating the angular momentum.

Thanks.